OpenStudy (anonymous):

An eleven pound projectile leaves the barrel of an experimental “super gun” with speed 9000 mph. The barrel is 155 feet long. Determine the acceleration rate, assuming it to be uniform.

OpenStudy (anitaasks):

I think this is how you do it (Final velocity)^2 = (Initial Velocity)^2 + 2ad Change feet into miles-> 155 feet (1mile/5280 feet) = (155/5280) (9000)^2 = 0 + 2(a)(155/5280) then solve for a.

OpenStudy (anonymous):

nope, still didnt work :( but thank you, I have to go to bed so you get best response for trying :)